# A lower bound for the first non-zero basic eigenvalue on a singular Riemannian foliation

### v2

@@ 5. The case of Positive Ricci curvature
We will follow the ideas that are showed in the work of Bakry-Qian (see Theorem 14, [BQ00]), and the work of Shi-Zhang [SZ07] to prove Theorem 1.3.
Lemma 5.1 ([BQ00]).
@@
To prove (1), see Theorem 13 of [BQ00].
For proving (2), see the discussion of eigenvalues on symmetric intervals of section 3 of [BQ00], or the last argument in Theorem 14 of [BQ00]. ∎
In the following lemma, we proceed exactly as in the proof of Theorem 14, [BQ00], with the new thing is we choose a more specific geodesic to take an integral, similar to the way that is introduced in the proof of Theorem 1.1.
@@
Use a similar argument as in the Theorem 14 of [BQ00]: First, let $u$ be the (basic) eigenfunction with the eigenvalue $\lambda_{1}^{B}$. Similar as the argument in the introduction of the subsection 4.1, we can normalize $u$ and assume $u$ is the (basic) eigenfunction with the eigenvalue $\lambda_{1}^{B}$ such that $\min u=-1$ and $0<\rho=\max u\leq 1$.
There exists an interval $[a,b]$ which has $\lambda=\lambda_{1}^{B}$ as the eigenvalue with the eigenfunction $v$ of $L_{K,n}v=-\lambda v$ with Neumann boundary conditions $v^{\prime}(a)=v^{\prime}(b)=0$, such that the corresponding eigenvector $v$ of the model satisfies v(a)=1v(a)=-1 and $v(b)=\rho$ (See Proposition 1, Section 3 and Theorem 9, Section 6, of [BQ00] for more detail).
Then, apply Theorem 8 of [BQ00] for the function $g=v^{-1}\circ u$ to get that $|\nabla g|\leq 1$.
@@
Recall the most important calculation from Shi-Zhang (see Lemma 2.2, [SZ07])
Lemma 5.3 (Lemma 2.2 of [SZ07]).
@@
Remarks.  When $\mathcal{F}$ is given as the (connected) fibers of a Riemannian submersion $\pi:M\to M/\mathcal{F}$ with totally geodesic (or, more generally, minimal) fibers, the basic spectrum of $M$ is turned out to be the spectrum of the base $M/\mathcal{F}$ of the submersion and $\lambda_{1}^{B}$ is the first non-zero eigenvalue $\alpha_{1}$ of the base $M/\mathcal{F}$. We see that if we apply directly Shi-Zhang estimate for the base $M/\mathcal{F}$ with (n1)K>0(n-1)K>0 as a lower bound for the Ricci curvature, we get a weaker estimate

### v3

@@ 5. The case of Positive Ricci curvature
We will follow the ideas that are showed in the work of Bakry-Qian (see Theorem 14, [3]), and the work of Shi-Zhang [25] to prove Theorem 1.3.
Lemma 5.1 ([3]).
@@
To prove (1), see Theorem 13 of [3].
For proving (2), see the discussion of eigenvalues on symmetric intervals of section 3 of [3], or the last argument in Theorem 14 of [3]. ∎
In the following lemma, we proceed exactly as in the proof of Theorem 14, [3], with the new thing is we choose a more specific geodesic to take an integral, similar to the way that is introduced in the proof of Theorem 1.1.
@@
Use a similar argument as in the Theorem 14 of [3]: First, let $u$ be the (basic) eigenfunction with the eigenvalue $\lambda_{1}^{B}$. Similar as the argument in the introduction of the subsection 4.1, we can normalize $u$ and assume $u$ is the (basic) eigenfunction with the eigenvalue $\lambda_{1}^{B}$ such that $\min u=-1$ and $0<\rho=\max u\leq 1$.
There exists an interval $[a,b]$ which has $\lambda=\lambda_{1}^{B}$ as the eigenvalue with the eigenfunction $v$ of $L_{K,n}v=-\lambda v$ with Neumann boundary conditions $v^{\prime}(a)=v^{\prime}(b)=0$, such that the corresponding eigenvector $v$ of the model satisfies v(a)=1v(a)=-1 and $v(b)=\rho$ (See Proposition 1, Section 3 and Theorem 9, Section 6, of [3] for more detail).
Then, apply Theorem 8 of [3] for the function $g=v^{-1}\circ u$ to get that $|\nabla g|\leq 1$.
@@
Recall the most important calculation from Shi-Zhang (see Lemma 2.2, [25])
Lemma 5.3 (Lemma 2.2 of [25]).
@@
Remark 5.1.
Theorem 1.3 can be applied particularly for $s=\frac{1}{2}$ to obtain Li’s type estimate:
(5.5) $\lambda_{1}^{B}\geq\frac{\pi^{2}}{\mathrm{d}_{M/\mathcal{F}}^{2}}+\frac{1}{2}(n-1)K$
We can actually learn more about which $s$ would give the best estimate: let $A=\frac{\pi^{2}}{\mathrm{d}_{M/\mathcal{F}}^{2}},B=(n-1)K$, then observe that the right hand side of the inequality 1.3 is a quadratic function of $s$:
$b(s)=4s(1-s)A+sB=-4As^{2}+(4A+B)s$
The only zero of $b(s)$ is $s=s_{0}=\frac{4A+B}{8A}$.
Therefore, by studying the function $b(s)$ on $(0,1)$, we can see that when $n\geq 3$:
  1. (1)
    If $s_{0}<1$, which is equivalent to
    $\mathrm{d}_{M/\mathcal{F}}<\frac{2}{\sqrt{(n-1)}}\frac{\pi}{\sqrt{K}},$
    then the maximum of $b(s)$ is
    $b(s_{0})=\frac{(4A+B)^{2}}{16A}=A+\frac{B}{2}+\frac{B^{2}}{16A}=\frac{\pi^{2}}{\mathrm{d}^{2}_{M/\mathcal{F}}}+\frac{1}{2}(n-1)K+\frac{B^{2}}{16A},$
    and it is sharper than the Li’s type estimate 5.5.
  2. (2)
    If $s_{0}\geq 1$, which is equivalent to say
    $\frac{2}{\sqrt{n-1}}\frac{\pi}{\sqrt{K}}\leq\mathrm{d}_{M/\mathcal{F}}\leq\frac{\pi}{\sqrt{K}}$
    (We recall the upper bound by Bonnet-Myer’s Theorem), then the supremum of $b(s)$ is by letting $s\to 1^{-}$, and we get $\lambda_{1}^{B}\geq(n-1)K$, which is weaker than Lichnerowicz’s estimate.
So, we can see that Theorem 1.3 is a sharp estimate when considering a small diameter of the leaf space, or even better, a small diameter of $M$. In this case, the optimal one is
$\lambda_{1}^{B}\geq\frac{\pi^{2}}{\mathrm{d}_{M/\mathcal{F}}^{2}}+\frac{1}{2}(n-1)K+\frac{(n-1)^{2}K^{2}\mathrm{d}_{M/\mathcal{F}}^{2}}{16\pi^{2}}.$
Remark 5.2.
When $\mathcal{F}$ is given as the (connected) fibers of a Riemannian submersion $\pi:M\to M/\mathcal{F}$ with totally geodesic (or, more generally, minimal) fibers, the basic spectrum of $M$ is turned out to be the spectrum of the base $M/\mathcal{F}$ of the submersion and $\lambda_{1}^{B}$ is the first non-zero eigenvalue $\alpha_{1}$ of the base $M/\mathcal{F}$. We see that if we apply directly Shi-Zhang estimate for the base $M/\mathcal{F}$ with (n1)K>0(n-1)K>0 as a lower bound for the Ricci curvature, we get a weaker estimate